OpenStudy (anonymous):

Which equation would you use to solve the following word problem? Sasha's mom decided to get her a cell phone for her birthday. Cell Plus has a plan that costs $22.95 per month plus an additional $0.08 per minute. Cell Best has a plan that costs $27.95 per month plus an additional $0.06 per minute. How many minutes can Sasha talk and have the same cost each month? 0.14m = 5.00 22.95 + 0.08m = 27.95 + 0.06m 22.95 + 0.08m = 28.01 27.95 + 0.06m = 23.03

OpenStudy (anonymous):

22.95 + 0.08m = 27.95 + 0.06m 0.08m - 0.06m = 27.95 - 22.95 0.02m = 5 m = 5/0.02 m = 250minutes